If each side of a square patio is increased by 4 feet, the area of the patio would be 196 square feet. Solve the equation (s +4)2 196 for s to find the length of a side of the patio. Provide your answer below: Sfeet ??
Answer #1

\large \left(s+4\right)^2=196

\large s+4=\sqrt{196}

\large s+4=\sqrt{14^2}}

\large s+4=14.................taking only positive sign because length is never negative

\large s=14-4

\large {\color{Red} s=10} feet

.

.